# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ A-D Summary of the Proof
From the above proof, we find that whether $\Psi(\gamma)$ is bounded or $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, Eq. (81) is enough to ensure that Eq. (22) holds with a given $\epsilon$. We let
@@
Therefore, to ensure that Eq. (81) holds with a given $\epsilon$, we can choose a $\Psi(\gamma)$ satisfying
${\color[rgb]{0,0,0}\Psi(\gamma)>\varsigma_{1}.}$ (89)
To ensure the validity of Eq. (89), we have to prove that $\varsigma_{1}$ is bounded. It is equivalent to proving that $\left(\frac{2K(x)}{V^{\frac{3}{2}}(x)}\right)^{2}$ is bounded on $[0,+\infty)$. For $\left(\frac{2K(x)}{V^{\frac{3}{2}}(x)}\right)^{2}$, its value is bounded for $x\to+\infty$ and finite $x$ as we analyzed above. For $x\to 0$, we have
@@
$=$ $\frac{2c_{0}}{(\log_{2}e)^{3}}\lim_{x\to 0}\frac{x^{\frac{3}{2}}}{(x+2)^{\frac{3}{2}}}\left(u_{4}x^{-\frac{3}{2}}+o\left(x^{-\frac{3}{2}}\right)\right)=u_{5},$
where $u_{4}$ and $u_{5}$ are finite constant. $o\left(x^{-\frac{3}{2}}\right)$ satisfies $\lim\limits_{x\to 0}\frac{o\left(x^{-\frac{3}{2}}\right)}{x^{-\frac{3}{2}}}=0$. Thus, we prove that $\varsigma_{1}$ is bounded. The proof is completed.
@@ Appendix B Proof of Theorem 3
For finite $|h|^{2}\gamma$, the proof is the same as that in Appendix A, in which $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ and $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$ are all bounded. Thus, in the following proof, we mainly focus on the analysis for $|h|^{2}\gamma\to+\infty$.
Note that we use the upper and lower bounds in the following proof. Thus, this result is under the assumption that
@@
Since $\epsilon(\gamma)\in(0,0.5]$, we have $1-\epsilon(\gamma)\geq\epsilon(\gamma)$. Thus, Eq. (91) is equivalent to
${\color[rgb]{0,0,0}\Psi(\gamma)>\left(\frac{2K(|h|^{2}\gamma)}{\epsilon(\gamma)V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2}.}$ (92)
Let $\varsigma_{2}=4\max_{x\geq 0}\left\{\frac{K^{2}(x)}{V^{3}(x)}\right\}$. Similar to Eq. (90), $\varsigma_{2}$ is bounded. Thus, we can let $\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2}$ to ensure Eq. (91) is satisfied. We let $\Psi(\gamma)\epsilon^{2}(\gamma)=(1+\omega)\varsigma_{2}$, where $\omega>0$ is an arbitrarily positive constant. Then we have
${\color[rgb]{0,0,0}\epsilon(\gamma)=\sqrt{\frac{(1+\omega)\varsigma_{2}}{\Psi(\gamma)}}.}$ (93)
Eq. (93) will be used below.
@@
Given that $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$, we have $1-\epsilon(\gamma)\to 1$ for $\gamma\to+\infty$. Thus, according to Eq. (32), $c_{1}\to+\infty$ for $\gamma\to+\infty$ since $\frac{dQ^{-1}(y)}{dy}\Big{|}_{y=1-\epsilon(\gamma)}\to-\infty$. Except for $c_{1}$, the other terms in $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ remains bounded. Thus, to ensure that there exists a finite $\nu$ such that $\lim\limits_{\gamma\to+\infty}\left|\frac{g_{u}(|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\right|\leq\nu$ holds, it is equivalent to requiring
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{\frac{dQ^{-1}(1-\epsilon(\gamma))}{dy}}{\Psi(\gamma)}<+\infty.}$ (94)
By substituting the derivative of inverse Gaussian Q-function $\frac{{\rm d}Q^{-1}(y)}{{\rm d}y}=\left[-\frac{1}{\sqrt{2\pi}}\exp\left(-\left(\frac{Q^{-1}(y)}{2}\right)^{2}\right)\right]^{-1}$ into Eq. (94), it is equivalent to requiring
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{e^{\left(\frac{Q^{-1}(1-\epsilon(\gamma))}{2}\right)^{2}}}{\Psi(\gamma)}=\lim_{\gamma\to+\infty}\frac{e^{\left(\frac{Q^{-1}(\epsilon(\gamma))}{2}\right)^{2}}}{\Psi(\gamma)}<+\infty.}$ (95)
Given that $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$ and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)\to+\infty$, the two terms in Eq. (80) are not bounded, which are the same as those mentioned in Appendix A. For $-\frac{1}{2}\log_{2}\Psi(\gamma)$, it follows Eq. (86) without changes. For simplicity, we let
${\color[rgb]{0,0,0}H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))=\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\right]}$ (96)
denote another unbounded term. Since $Q^{-1}(x)$ is a decreasing function and $\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}>0$, we have
${\color[rgb]{0,0,0}Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\leq Q^{-1}\left(1-\epsilon(\gamma)\right)+Q^{-1}(\epsilon(\gamma))=0.}$ (97)
Based on Eq. (97), we have
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\leq 0.}$ (98)
We have proved the upper bound of $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}$ is finite. Then, we demonstrate the condition with which the lower bound of $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}$ is finite. According to Eq. (93) and the definition of $\varsigma_{2}$, we have
@@
We will give the proof of $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}=0$ in Appendix C. Thus, $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}<+\infty$ holds if
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{\Psi(\gamma)}}<+\infty.}$ (101)
Thus, if Eq. (101) holds, we have $\lim\limits_{\gamma\to+\infty}\left|\frac{g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\right|\leq\nu$.
Therefore, if Eqs. (95), (101), and $\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2}$ hold, the condition shown in Eq. (22) is satisfied. The proof is completed.
Since $\Psi(\gamma)\epsilon^{2}(\gamma)=(1+\omega)\varsigma_{2}$, we have $\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)<\epsilon(\gamma)<\frac{1}{2}$. Thus, we obtain $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}\geq 0$. To adopt the squeeze theorem, we then need to demonstrate that $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}\leq 0$. For simplicity, we let $\omega_{1}=\sqrt{\varsigma_{2}}\left(\sqrt{1+\omega}-1\right)$ in the following proof. We let $Q^{-1}\left(v\right)=x$, which can also be written as $v=Q(x)$. Based on Mills’ ratio, for $x>0$, we obtain that
${\color[rgb]{0,0,0}\frac{1}{\sqrt{2\pi}}\frac{x}{1+x^{2}}e^{-\frac{x^{2}}{2}}\leq v\leq\frac{1}{\sqrt{2\pi}}\frac{1}{x}e^{-\frac{x^{2}}{2}}.}$ (102)
Then, we have
${\color[rgb]{0,0,0}x^{2}+\ln x^{2}\leq-2\ln\left(\sqrt{2\pi}v\right).}$ (103)
${\color[rgb]{0,0,0}-\sqrt{W\left(\frac{v^{-2}}{2\pi}\right)}\leq x\leq\sqrt{W\left(\frac{v^{-2}}{2\pi}\right)}.}$ (104)
Since $\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}<\frac{1}{2}$, we have
${\color[rgb]{0,0,0}0<Q^{-1}\left(\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}\right)<\sqrt{W\left(\frac{\Psi(\gamma)}{2\pi\omega_{1}^{2}}\right)}.}$ (105)
Based on Eq. (105), we obtain
@@
Eq. (106a) holds because, according to [48], for $y>e$, the Lambert W function satisfies
${\color[rgb]{0,0,0}\ln y-\ln\ln y+\frac{\ln\ln y}{2\ln y}\leq W(y)\leq\ln y-\ln\ln y+\frac{e}{e-1}\frac{\ln\ln y}{\ln y}.}$ (107)
Thus, for $y>e$, we obtain
${\color[rgb]{0,0,0}\lim_{y\to+\infty}\frac{\sqrt{W(y)}}{\sqrt{\log_{2}y}}=\frac{1}{\sqrt{\log_{2}e}}.}$ (108)
@@ Appendix D Proof of Theorem 5
Based on the definition of $s_{\infty}$ in Eq. (44) and $\tilde{R}^{\epsilon}_{N}[n]=R^{\epsilon}_{N}[n]-G(N,|h[n]|^{2}\gamma)$, we have
@@
Since $N=\Psi(\gamma)$ and $N\theta=\varrho$, we have $\theta=\frac{\varrho}{\Psi(\gamma)}$. In Section II, we have proved that if $\Psi(\gamma)$ satisfies Theorem 2, Eq. (22) holds with a given $\epsilon$. Thus, we have
${\color[rgb]{0,0,0}\left|\theta G(N,|h[n]|^{2}\gamma)\right|=\left|\frac{\varrho G(N,|h[n]|^{2}\gamma)}{\Psi(\gamma)}\right|\leq\varrho\nu.}$ (110)
Thus, we have that $c_{5}\leq e^{-T\theta G(N,|h[n]|^{2}\gamma)}\leq c_{6}$, where $c_{i}$ is a positive constant, $i=5,6$. Thus, under the above conditions, we obtain
According to Eq. (111), it is obvious that $s_{\infty}=1$ for the AWGN channel. For the Nakagami-$m$ fading channel, we have
@@
We then divide the proof into four parts. First, we focus on the upper and lower bounds of $s_{\infty}$ for $m\geq 1$. Then we present the upper and lower bounds of $s_{\infty}$ for $m\in[0.5,1)$. For simplicity, we let $\beta=\sqrt{N}Q^{-1}(\epsilon)\log_{2}e$ in this paper.
D-A Upper bound of $s_{\infty}$ for $m\geq 1$
For $m\geq 1$, we first derive the upper bound of $s_{\infty}$. Since $R_{\epsilon}[n]\leq N\log_{2}\left(1+|h[n]|^{2}\gamma\right)$, we obtain
@@ D-A1 $1<\frac{m}{\theta NT\log_{2}e}$
According to [14], $0<\int_{0}^{\infty}z^{a}e^{-x}{\rm d}x<\infty$ if $a>-1$. Since in this case $m>\theta NT\log_{2}e$, we obtain $m-\theta NT\log_{2}e-1>-1$. Therefore, $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ is bounded by a finite constant. Thus, under the condition $1<\frac{m}{\theta NT\log_{2}e}$, Eq. (113) can be further expressed as
$s_{\infty}\leq\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}=1.$ (115)
Eq. (115) holds because $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}N}{N}=\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\Psi(\gamma)}{2\Psi(\gamma)}=0$ holds under the condition that $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$.
D-A2 $1>\frac{m}{\theta NT\log_{2}e}$
For $1>\frac{m}{\theta NT\log_{2}e}$, we focus on the integral $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$. We find that

### v4

@@ A-D Summary of the Proof
From the above proof, we find that whether $\Psi(\gamma)$ is bounded or $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, Eq. (82) is sufficient to ensure that Eq. (22) holds for a given $\epsilon$. For the convenience of expression, we define
@@
Therefore, to ensure that Eq. (82) holds for a given $\epsilon$, we can choose a $\Psi(\gamma)$ satisfying
$\Psi(\gamma)>\varsigma_{1}.$ (90)
To establish the validity of Eq. (90), we have to prove that $\varsigma_{1}$ is bounded. It is equivalent to proving that $\left(\frac{2K(x)}{V^{\frac{3}{2}}(x)}\right)^{2}$ is bounded on the interval $[0,+\infty)$. For $\left(\frac{2K(x)}{V^{\frac{3}{2}}(x)}\right)^{2}$, we have shown that its value is bounded for both $x\to+\infty$ and finite $x$. As $x\to 0$, we have
@@
$=$ $\frac{2c_{0}}{(\log_{2}e)^{3}}\lim_{x\to 0}\frac{x^{\frac{3}{2}}}{(x+2)^{\frac{3}{2}}}\left(u_{4}x^{-\frac{3}{2}}+o\left(x^{-\frac{3}{2}}\right)\right)$
$=$ $u_{5},$
where $u_{4}$ and $u_{5}$ are finite constant, and $o\left(x^{-\frac{3}{2}}\right)$ satisfies $\lim\limits_{x\to 0}\frac{o\left(x^{-\frac{3}{2}}\right)}{x^{-\frac{3}{2}}}=0$. Thus, we have proved that $\varsigma_{1}$ is bounded. The proof is completed.
@@ Appendix B Proof of Theorem 3
For finite $|h|^{2}\gamma$, the proof is the same as that in Appendix A, where both $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ and $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$ are proved to be bounded. Therefore, in the following proof, we mainly focus on the analysis for $|h|^{2}\gamma\to+\infty$.
Note that we use the upper and lower bounds in the following proof. Therefore, this result is based on the assumption shown in Eq. (92), which is located at the bottom of this page.
@@
$H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))=\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\right].$ (97)
$Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\leq Q^{-1}\left(1-\epsilon(\gamma)\right)+Q^{-1}(\epsilon(\gamma))=0.$ (98)
Since $\epsilon(\gamma)\in(0,0.5]$, it follows that $1-\epsilon(\gamma)\geq\epsilon(\gamma)$. Thus, to ensure Eq. (92) holds, it is sufficient to ensure the following condition:
$\Psi(\gamma)>\left(\frac{2K(|h|^{2}\gamma)}{\epsilon(\gamma)V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2}.$ (93)
We define
$\varsigma_{2}=4\max_{x\geq 0}\left\{\frac{K^{2}(x)}{V^{3}(x)}\right\}.$
Similar to the analysis in Eq. (91), $\varsigma_{2}$ is bounded. Thus, we can impose the condition $\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2}$ to ensure Eq. (92) is satisfied. We let $\Psi(\gamma)\epsilon^{2}(\gamma)=(1+\omega)\varsigma_{2}$, where $\omega>0$ is an arbitrarily positive constant. Then we have
$\epsilon(\gamma)=\sqrt{\frac{(1+\omega)\varsigma_{2}}{\Psi(\gamma)}}.$ (94)
Eq. (94) will be used in the subsequent analysis.
@@
Given that $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$, we have $1-\epsilon(\gamma)\to 1$ as $\gamma\to+\infty$. Thus, according to Eq. (33), $c_{1}\to+\infty$ for $\gamma\to+\infty$ since $\frac{dQ^{-1}(y)}{dy}\Big{|}_{y=1-\epsilon(\gamma)}\to-\infty$. Except for $c_{1}$, the other terms in $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ remain bounded. Therefore, to ensure that there exists a finite $\nu$ such that $\lim\limits_{\gamma\to+\infty}\left|\frac{g_{u}(|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\right|\leq\nu$ holds, it is equivalent to requiring
$\lim_{\gamma\to+\infty}\frac{\frac{dQ^{-1}(1-\epsilon(\gamma))}{dy}}{\Psi(\gamma)}<+\infty.$ (95)
By substituting the derivative of inverse Gaussian Q-function, which is given by
$\frac{{\rm d}Q^{-1}(y)}{{\rm d}y}=\left[-\frac{1}{\sqrt{2\pi}}\exp\left(-\left(\frac{Q^{-1}(y)}{2}\right)^{2}\right)\right]^{-1},$
into Eq. (95), we arrive at the requirement that
$\lim_{\gamma\to+\infty}\frac{e^{\left(\frac{Q^{-1}(1-\epsilon(\gamma))}{2}\right)^{2}}}{\Psi(\gamma)}$ (96)
$=$ $\lim_{\gamma\to+\infty}\frac{e^{\left(\frac{Q^{-1}(\epsilon(\gamma))}{2}\right)^{2}}}{\Psi(\gamma)}<+\infty.$
@@
Given that $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$ and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)\to+\infty$, the two terms in Eq. (81) are unbounded, which are the same as those mentioned in Appendix A. The term $-\frac{1}{2}\log_{2}\Psi(\gamma)$ follows Eq. (87) without changes. For simplicity, we define $H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))$ in Eq. (97), which is located at the bottom of this page, to represent another unbounded term. Since $Q^{-1}(x)$ is a decreasing function and $\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}>0$, we obtain Eq. (98), which is also located at the bottom of this page. From Eq. (98), we have
$\lim_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\leq 0.$ (99)
We have proved that the upper bound of $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}$ is finite. Next, we demonstrate the condition under which the lower bound of $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}$ is finite. According to Eq. (94) and the definition of $\varsigma_{2}$, we derive Eq. (100) at the bottom of the next page. Based on Eq. (100), we obtain Eq. (101), which is also located at the bottom of the next page.
We will provide the proof of $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}=0$ in Appendix C. Thus, $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}<+\infty$ holds if
$\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{\Psi(\gamma)}}<+\infty.$ (102)
Thus, if Eq. (102) holds, we have
$\lim\limits_{\gamma\to+\infty}\left|\frac{g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\right|\leq\nu.$
Therefore, if Eqs. (96), (102), and $\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2}$ hold, the condition stated in Eq. (22) is satisfied. The proof is completed.
Since $\Psi(\gamma)\epsilon^{2}(\gamma)=(1+\omega)\varsigma_{2}$, we have $\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)<\epsilon(\gamma)<\frac{1}{2}$. Thus, we obtain $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}\geq 0$. To apply the squeeze theorem, we then need to demonstrate that $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}\leq 0$. For simplicity, we define $\omega_{1}=\sqrt{\varsigma_{2}}\left(\sqrt{1+\omega}-1\right)$ in the following proof. Let $Q^{-1}\left(v\right)=x$, which implies $v=Q(x)$. Based on Mills’ ratio, for $x>0$, we obtain that
$\frac{1}{\sqrt{2\pi}}\frac{x}{1+x^{2}}e^{-\frac{x^{2}}{2}}\leq v\leq\frac{1}{\sqrt{2\pi}}\frac{1}{x}e^{-\frac{x^{2}}{2}}.$ (103)
Then, we have
$x^{2}+\ln x^{2}\leq-2\ln\left(\sqrt{2\pi}v\right).$ (104)
$-\sqrt{W\left(\frac{v^{-2}}{2\pi}\right)}\leq x\leq\sqrt{W\left(\frac{v^{-2}}{2\pi}\right)}.$ (105)
Since $\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}<\frac{1}{2}$, we obtain the following relationships from Eq. (105):
$0<Q^{-1}\left(\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}\right)<\sqrt{W\left(\frac{\Psi(\gamma)}{2\pi\omega_{1}^{2}}\right)}.$ (106)
Based on Eq. (106), we obtain Eq. (107), which is presented at the bottom of this page. Eq. (107a) holds because, according to [31], for $y>e$, the Lambert W function satisfies the following inequalities:
$W(y)\geq\ln y-\ln\ln y+\frac{\ln\ln y}{2\ln y},$ (108)
and
$W(y)\leq\ln y-\ln\ln y+\frac{e}{e-1}\frac{\ln\ln y}{\ln y}.$ (109)
Based on Eqs. (108) and (109), we find that for $y>e$, the following equation holds:
$\lim_{y\to+\infty}\frac{\sqrt{W(y)}}{\sqrt{\log_{2}y}}=\frac{1}{\sqrt{\log_{2}e}}.$ (110)
@@ Appendix D Proof of Theorem 5
Based on the definition of $s_{\infty}$ proposed in Eq. (45) and $\tilde{R}^{\epsilon}_{N}[n]=R^{\epsilon}_{N}[n]-G(N,|h[n]|^{2}\gamma)$, we obtain Eq. (111) at the bottom of this page.
Since $N=\Psi(\gamma)$ and $N\theta=\varrho$, we have $\theta=\frac{\varrho}{\Psi(\gamma)}$. As discussed in Section II, if $\Psi(\gamma)$ satisfies the conditions of Theorem 2, Eq. (22) holds for a given $\epsilon$. Thus, the following inequality holds:
$\left|\theta G(N,|h[n]|^{2}\gamma)\right|=\left|\frac{\varrho G(N,|h[n]|^{2}\gamma)}{\Psi(\gamma)}\right|\leq\varrho\nu.$ (112)
From Eq. (112), we obtain that $c_{5}\leq e^{-T\theta G(N,|h[n]|^{2}\gamma)}\leq c_{6}$, where $c_{5}$ and $c_{6}$ are positive constants. Under the above conditions, we obtain
From Eq. (113), it is obvious that $s_{\infty}=1$ for the AWGN channel. For the Nakagami-$m$ fading channel, we obtain Eq. (114), which is presented at the bottom of this page.
We will utilize Eq. (114) to carry out the following proof. The following proof is divided into four parts. We analyze the upper bound of $s_{\infty}$ for $m\geq 1$ in the first part, followed by its lower bound in the second part. Next, we present the upper and lower bounds of $s_{\infty}$ for $m\in[0.5,1)$ in the third and fourth part, respectively. For simplicity, we let $\beta=\sqrt{N}Q^{-1}(\epsilon)\log_{2}e$ in the proofs of this paper.
D-A Upper bound of $s_{\infty}$ for $m\geq 1$
For $m\geq 1$, we first derive the upper bound of $s_{\infty}$. Since $R_{\epsilon}[n]\leq N\log_{2}\left(1+|h[n]|^{2}\gamma\right)$, we derive Eq. (115), which is presented at the bottom of this page.